# Cognitive Friction: A Decision-Theoretic Framework for Bounded Deliberation in Tool-Using Agents

### 01 Apr 2026 v2

@@ Abstract
Autonomous tool-using agents operating in networked environments must decide both which information source to query and when to stop querying and act. Without principled bounds on information-acquisition costs, unconstrained agents exhibit systematic failure modes: excessive tool use under congestion, prolonged deliberation under time decay, and brittle behavior under ambiguous evidence. We propose the Triadic Cognitive Architecture (TCA), a unified decision-theoretic framework that formalizes these failure modes through the concept of Cognitive Friction. By synthesizing nonlinear filtering theory, congestion-dependent cost dynamics, and HJB optimal stopping, we model deliberation as a stochastic control problem over a joint belief–congestion state space, where information acquisition is explicitly priced by tool-dependent signal quality and live network load. Rather than relying on arbitrary heuristic stop-tokens or fixed query budgets, TCA derives an HJB-inspired stopping boundary and instantiates a computable rollout-based approximation of belief-dependent value-of-information with a net-utility halting condition. We validate the framework on two controlled simulation environments, the Emergency Medical Diagnostic Grid (EMDG) and the Network Security Triage Grid (NSTG), designed to isolate the decision-theoretic quantities of interest (stopping quality, action selection under congestion, and temporal urgency) under reproducible conditions. TCA reduces time-to-action while improving resource outcomes without degrading accuracy: over greedy baselines, TCA gains 36 viability points in EMDG and 33 integrity points in NSTG. Two purposive stopping baselines confirm that the joint optimization of selection and stopping is essential; stopping rules alone recover at most 4 viability points. A univariate sensitivity sweep over $\alpha$, $\beta$, and $\lambda_{S}$ shows stable accuracy and interpretable tradeoffs, and an empirical sweep over $\eta\in\{0,0.1,0.3,0.5\}$ confirms $\eta{=}0$ is optimal on EMDG trajectories under high temporal urgency.
@@ 1 Introduction
Why this is new.
A skeptical reader might object: “This is just rational metareasoning russell1991right (stop when computation cost exceeds utility gain), or classical VOI-based sequential planning, or cost-sensitive bandit stopping with a fancy motivation.” The critical distinction is that these prior frameworks typically treat information-acquisition costs as static scalars fixed independently of the agent’s state. TCA instead makes costs dynamically state-dependent: the spatial friction $\lambda_{S}(C_{t}{+}\Omega(a))$ depends on the live congestion state $C_{t}$, and the temporal friction $\beta\,(t{+}\tau(a))$ is parameterized by the same rate constant $\beta$ that governs the exponential resource decay; both accumulate as deliberation proceeds. TCA jointly models (i) topology- and congestion-dependent information access costs (Space), (ii) time-dependent opportunity costs under delay (Time), and (iii) belief-dependent value-of-information (Truth), within a single stopping framework derived from optimal stopping theory, not ad hoc budget rules. The key contribution is coupling these three state-dependent terms in a unified decision rule and providing a computable rollout-based approximation to the resulting stopping boundary.
@@ 2.2 Bounded Rationality and Meta-Reasoning
The necessity of pricing deliberation traces back to Simon’s foundational concept of bounded rationality simon1955behavioral , which posits that agents must make decisions under the strict limits of available information, cognitive capacity, and time. This was formally extended into meta-reasoning and bounded optimality by Russell and Wefald russell1991right , who argued that an agent should compute only until the cost of computation outweighs the expected increase in decision utility. In deep learning, Adaptive Computation Time (ACT) graves2016adaptive introduced halting mechanisms for recurrent networks based on internal confidence. TCA extends these classical concepts to tool-using autonomous agents, formalizing sequential tool queries as a continuous-time physical process subject to network congestion and temporal decay.
@@ Results.
Table 7 reports terminal statistics. The fraction of (configuration, seed) pairs in which TCA selects a non-minimum-latency tool at step 0 rises monotonically: $76\%$ at $|\mathcal{A}|{=}5$, $85\%$ at $|\mathcal{A}|{=}10$, and $90\%$ at $|\mathcal{A}|{=}20$. (For comparison, this fraction is $0\%$ in the 2-action EMDG, where Hematology Lab is always preferred over MRI; see Table 1.) TCA terminal viability remains stable at 91.891.893.993.9 across all sizes, fully consistent with the main-text EMDG result of $93.03\pm 0.19$. TCA consistently outperforms ReAct (gap $\geq+9.7$ viability points; accuracy $1.00$ throughout).
The absolute gap is smaller than in the 2-action EMDG ($\approx$10–15 vs. 36 points), which is expected: in random tool configurations, the greedy max-gain rule occasionally selects a low-latency tool when the highest-gain tool happens to have moderate latency, reducing, but never eliminating, TCA’s advantage. Crucially, TCA dominates in every configuration tested and its viability is stable across $|\mathcal{A}|$, confirming that the controller scales gracefully and performs genuine cost-aware optimization rather than a degenerate “always pick cheapest” rule. We view this sweep as a robustness check that the policy does not collapse to trivial cost minimization; it is not intended as an exhaustive scaling benchmark.

### 02 Apr 2026 v3

@@ Abstract
Autonomous tool-using agents operating in networked environments must decide both which information source to query and when to stop querying and act. Without principled bounds on information-acquisition costs, unconstrained agents exhibit systematic failure modes: excessive tool use under congestion, prolonged deliberation under time decay, and brittle behavior under ambiguous evidence. We propose the Triadic Cognitive Architecture (TCA), a decision-theoretic framework that formalizes these failure modes through Cognitive Friction. By synthesizing nonlinear filtering theory, congestion-dependent cost dynamics, and HJB optimal stopping, we model deliberation as a stochastic control problem over a joint belief–congestion state space, where information acquisition is explicitly priced by tool-dependent signal quality and live network load. Rather than relying on heuristic stop-tokens or fixed query budgets, TCA yields an HJB-inspired stopping boundary and a computable rollout-based approximation of belief-dependent value-of-information with a net-utility halting condition. We validate TCA on two controlled simulation environments, the Emergency Medical Diagnostic Grid (EMDG) and the Network Security Triage Grid (NSTG), designed to isolate stopping quality, action selection under congestion, and temporal urgency under reproducible conditions. TCA reduces time-to-action while improving resource outcomes without degrading accuracy: over greedy baselines, TCA gains 36 viability points in EMDG and 33 integrity points in NSTG, and purposive stopping baselines confirm that stopping rules alone recover at most 4 viability points. Finally, we provide an illustrative instantiation around a black-box LLM on a memorisation-free corpus, showing that the same stopping principle executes using empirically computable uncertainty and VOI proxies.
@@ 2.2 Bounded Rationality and Meta-Reasoning
The necessity of pricing deliberation traces back to Simon’s foundational concept of bounded rationality simon1955behavioral , which posits that agents must make decisions under the strict limits of available information, cognitive capacity, and time. This was formally extended into meta-reasoning and bounded optimality by Russell and Wefald russell1991right , who argued that an agent should compute only until the cost of computation outweighs the expected increase in decision utility. In deep learning, Adaptive Computation Time (ACT) graves2016adaptive introduced halting mechanisms for recurrent networks based on internal confidence. TCA extends these classical concepts to tool-using autonomous agents, formalizing sequential tool queries as a continuous-time physical process subject to network congestion and temporal decay. Crucially, classical meta-reasoning and VOI-based planners typically treat information-acquisition costs as static scalars independent of agent state. TCA instead makes costs dynamically state-dependent: spatial friction grows with live congestion $C_{t}$ and temporal friction accumulates with delay $t$, both changing as deliberation proceeds. This coupling of state-dependent cost dynamics with belief-dependent VOI within a single HJB stopping framework is the key structural departure from prior bounded rationality work.
@@ Results.
Table 7 reports terminal statistics. The fraction of (configuration, seed) pairs in which TCA selects a non-minimum-latency tool at step 0 rises monotonically: $76\%$ at $|\mathcal{A}|{=}5$, $85\%$ at $|\mathcal{A}|{=}10$, and $90\%$ at $|\mathcal{A}|{=}20$. (For comparison, this fraction is $0\%$ in the 2-action EMDG, where Hematology Lab is always preferred over MRI; see Table 1.) TCA terminal viability remains stable at 91.891.893.993.9 across all sizes, fully consistent with the main-text EMDG result of $93.03\pm 0.19$. TCA consistently outperforms ReAct across all sizes tested (gaps: +9.8+9.8 at $|\mathcal{A}|{=}5$, +9.7+9.7 at $|\mathcal{A}|{=}10$, +15.0+15.0 at $|\mathcal{A}|{=}20$; accuracy $1.00$ throughout). Notably, the gap at $|\mathcal{A}|{=}20$ is larger than at $|\mathcal{A}|{=}5$: with more tools, the greedy max-gain baseline encounters misaligned cost-quality tradeoffs more frequently, while TCA’s joint optimization handles them correctly. In this sweep, TCA’s advantage is not eroded by larger action spaces.
The absolute gap is smaller than in the 2-action EMDG ($\approx$10–15 vs. 36 points), which is expected: in random tool configurations, the greedy max-gain rule occasionally selects a low-latency tool when the highest-gain tool happens to have moderate latency, reducing but never eliminating TCA’s advantage. TCA dominates in every configuration tested and its viability is stable across $|\mathcal{A}|$, confirming that the controller scales gracefully. We view this sweep as a robustness check; it is not intended as an exhaustive scaling benchmark.
@@ Appendix C Discrete Controller Pseudocode
Appendix D Real-LLM Instantiation on a Controlled Corpus
The main experiments (Sections 44.9) use a synthetic simulation to isolate belief and cost dynamics under controlled conditions. We complement this with an illustrative instantiation on a real LLM to confirm that the TCA stopping principle is directly implementable around a black-box language model.
Experimental disclaimer.
These experiments are illustrative. The primary contribution of TCA concerns decision-theoretic stopping under observable belief and cost, which we validate in the controlled EMDG/NSTG environments. The real-LLM instantiation shows that the controller is executable without modification when belief and VOI are replaced by empirically computable proxies.
Setup.
We construct a fictional closed-world corpus of 25 short documents describing invented nations, persons, treaties, and scientific facts (e.g., the Porvine Confederation, element Zethanite). Fictional content is essential: it guarantees that GPT-4.1 has zero memorised knowledge of the corpus, so any accuracy difference between policies reflects the controller logic rather than parametric recall. We pair the corpus with 60 short-answer questions whose answers appear in exactly one document.
Belief proxy.
At each decision point we call GPT-4.1 N=5N{=}5 times at temperature T=0.7T{=}0.7 with the prompt below and compute an empirical entropy $H_{\text{pre}}$ over normalised answers, used as a proxy for belief uncertainty. This self-consistency entropy is labelled as a proxy throughout.
System: You answer factual questions with a short, precise phrase or number. Output only the answer, no explanation.
User: Question: {q}. Answer with a short phrase or number.
VOI proxy.
We estimate $\widehat{\Delta H}$ via a one-step cloned rollout: retrieve the top-3 documents by keyword overlap (no external embedding or retriever), form a context string, re-sample N=5N{=}5 answers, and compute $\widehat{\Delta H}=H_{\text{pre}}-H_{\text{post}}$. We denote this proxy entropy by $H_{\text{pre}}$ before retrieval and $H_{\text{post}}$ after retrieval (both computed from N=5N{=}5 samples); these are proxy quantities distinct from the categorical belief entropy in the EMDG/NSTG experiments, which is computed from an explicit probability vector. We deliberately use a weak retriever (keyword overlap, no embeddings) to avoid confounding the stopping behaviour with retriever quality. This mirrors the rollout-based VOI approximation used in EMDG (Section 3.4).
Stop rule.
The TCA stop rule is applied without modification:
$\text{retrieve iff}\quad\widehat{\Delta H}-\alpha\,c_{r}>0,\qquad\alpha{=}1.0,\quad c_{r}{=}1.0.$
Let $H_{\text{final}}=H_{\text{post}}$ if retrieval is taken, else $H_{\text{pre}}$. After stopping: Act if $H_{\text{final}}\leq\varepsilon_{1}{=}0.25$, else Defer.
Baselines.
NeverRetrieve: answer immediately from parametric knowledge alone. AlwaysRetrieve: retrieve once unconditionally, then answer. TCA: retrieve only when net utility is positive.
Utility.
$U=+10\cdot\mathbf{1}[\text{correct}]-10\cdot\mathbf{1}[\text{wrong}]-c_{r}\cdot n_{\text{retrieve}}$, matching the paper’s cost-reward structure.
Results.
Table 9: Real-LLM instantiation on fictional corpus (N=60N{=}60 questions, GPT-4.1). $U$ = expected utility; EM = exact match; $n_{r}$ = avg retrieval calls; %ACT, %DEFER = terminal rates. NeverRetrieve scores EM = 0.00 confirming zero memorisation contamination. TCA uses 63% of AlwaysRetrieve’s retrieval budget while exhibiting both terminal outcomes (Act/Defer) under the same net-utility stopping rule.
Policy $U$ EM F1 $n_{r}$ %ACT %DEFER
NeverRetrieve 10.00-10.00 $0.000$ $0.054$ $0.00$ $100\%$ $0\%$
AlwaysRetrieve +5.33+5.33 $0.817$ $0.927$ $1.00$ $100\%$ $0\%$
TCA (ours) +2.20+2.20 $0.500$ $0.580$ $0.63$ $72\%$ $28\%$
Analysis.
This appendix is a validation experiment, not a performance benchmark. Its purpose is to confirm that the TCA controller, when instantiated with a real LLM on a memorisation-free corpus, produces the full behavioural structure predicted by the theory: selective retrieval and two terminal outcomes (Act/Defer). The retrieval cost $c_{r}{=}1$ is set to be non-trivial so that the Defer path is exercised: for a subset of questions the estimated VOI does not justify retrieval, whereas for others retrieval remains beneficial (hence AlwaysRetrieve can still achieve higher utility in this regime). With $c_{r}{\approx}0$, TCA trivially degenerates to AlwaysRetrieve and Defer is never triggered, making the experiment uninformative about the stopping rule. At this cost level the controller correctly defers on 28% of questions and acts on 72%, exhibiting selective retrieval and both terminal behaviours in a single run.
Of the 38 Act+retrieve episodes, 79% are correct. The 5 Act-without-retrieve episodes ($H_{\text{pre}}{=}0$; all five samples identical at T=0.7T{=}0.7) illustrate a known limitation of self-consistency as a belief proxy: when GPT-4.1 is parametrically overconfident on unknown facts the proxy underestimates true uncertainty. All 5 were incorrect. We disclose this as a fidelity limitation of the proxy, not a failure of the stopping rule itself.
TCA’s expected utility (+2.20) is lower than AlwaysRetrieve (+5.33): this is the expected consequence of evaluating TCA in a regime where retrieval has non-trivial cost and the controller sometimes prefers deferral. The utility comparison does not assess whether TCA is a better retrieval policy; it assesses whether the controller correctly identifies when retrieval is not worth its cost; 17 Defer decisions confirm that it does. NeverRetrieve’s EM = 0.000 confirms the corpus is memorisation-free.